• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 08:24:08 on November 02, 2008

    Entry number 246200

    keyword=Owl Shift Summary

    Owl Shift Summary

    12:00 Shift started while the MCC are trying to deliver us a good beam. The requested current was 5 uA beam (raster off). The optics (BeO) target is in.

    12:55 The beam is usable now with X and Y RMS values between 100 and 200 um.

    1:05 Taking a 2 uA beam (raster off) run for 10 min while setting the BB wire chamber parameters

    1:20 Changing the raster to 4x4 (MCC) and take a run for another 10 min.

    1:30 Moving the target to the reference cell (H2 at 100 psig)

    1:35 Asking for 4 uA beam with BPMA at -1.5,+2.4 and BPMB at -1.5,+2.5

    1:45 Changed the beam back to 2 uA while setting the BB wire chamber parameters.

    1:55 Start 1.5 h of about 30 minutes runs with T1 at 800 Hz.

    2:00 Clock changes from 2AM back to 1AM (the end of the Daylight Saving Time).

    2:15 Frequent bigbite wire chamber trips when the beam comes ramps up. Had to stop the run and start a new one every time.

    2:45 Started the checks of the wire chamber parameters. Taking 100k event runs while changing the HV to 1500, 1550, 1600, 1650V and the Threshold to 3.0, 3.5, 4.0, 4.5, 5.0V for each H V value. We could not do the 1700 and 1750V HV checks due to tripping of the BB wire chambers.

    4:45 Paged Kaylan to discuss the problem we are experiencing with ROC9.

    5:00 Beam is very trippy.

    5:30 Paged JP to discuss the effect of the unstable beam on the He3 Cell. He recommended to always make sure of the exact beam position before moving to cell.

    6:10 Paged the run coordinator to inform him about the ROC9 problem. He managed to reboot it remotely. We were able to take a couple of other runs before the beam is gone.

    7:00 MCC expect the beam will be away for 1 hour while checking the cavity problems.

    7:55 MCC called to inform us that there is a cavity problem which might take another 40 min. to fix.

    8:00 Shift ends.